# On the Statistical Complexity of Estimating VENDI Scores from Empirical Data

### v1

@@ A.5 Proof of Theorem 3
Proof of Part (a). As defined by [4], the FKEA method uses the eigenvalues of $t$ random Fourier frequencies $\omega_{1},\ldots,\omega_{t}$ where for each $\omega_{i}$ they consider two features $\cos(\omega_{i}^{\top}x)$ and $\sin(\omega_{i}^{\top}x)$. Following the definitions, it can be seen that $k(x,x^{\prime})=\mathbb{E}_{\omega\sim p_{\omega}}\bigl{[}\cos(\omega^{\top}(x-x^{\prime}))\bigr{]}$ which is approximated by FKEA as $\frac{1}{t}\sum_{i=1}^{t}\cos(\omega_{i}^{\top}(x-x^{\prime}))$. Therefore, if we define kernel matrix $K_{i}$ as the kernel matrix for $k_{i}(x,x^{\prime})=\cos(\omega_{i}^{\top}(x-x^{\prime}))$, then we will have
@@
On the other hand, we note that $\|\frac{1}{n}K\|_{F}\leq 1$ holds as the kernel function is normalized and hence $|k(x,x^{\prime})|\leq 1$. Since the Frobenius norm is the $\ell_{2}$-norm of the vectorized version of the matrix, we can apply Vector Bernstein inequality in Lemma 1 to show that for every $0\leq\epsilon\leq 2$:
@@
Then, we apply the Hoffman-Wielandt inequality to show that for the sorted eigenvalue vectors of $\frac{1}{n}K$ (denoted by $\widehat{\boldsymbol{\lambda}}_{n}$) and $\frac{1}{n}K^{\mathrm{FKEA}(t)}$ (denoted by ${\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}$) we will have $\|\widehat{\boldsymbol{\lambda}}_{n}-{\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}\|_{2}\leq\|\frac{1}{n}K^{\mathrm{FKEA}(t)}-\frac{1}{n}K\|_{F}$, which together with the previous inequality leads to
@@
Proof of Part (b). To show this theorem, we use Theorem 3 from [30], which shows that if the $r$th largest eigenvalue of the kernel matrix $\frac{1}{n}K$ satisfies $\lambda_{r}\leq\frac{\tau}{n}$, then given $t\geq Cr\log(n)$ ($C$ is a universal constant), the following spectral norm bound will hold with probability $1-\frac{2}{n^{3}}$:
@@ Appendix B Additional Numerical Results
In this section, we present supplementary results concerning the evaluation of diversity and the convergence behavior of different variants of the VENDI score. We extend the convergence experiments discussed in the main text to include the truncated StyleGAN3-t FFHQ dataset (Figure LABEL:VENDI_stylegan3_convergence) and the StyleGAN-XL ImageNet dataset (Figure LABEL:VENDI_styleganxl_convergence). Furthermore, we demonstrate that the truncated VENDI statistic effectively captures the diversity characteristics across various data modalities. Specifically, we conducted similar experiments as shown in Figures LABEL:VENDI_imagenet_truncation and LABEL:VENDI_ffhq_truncation on text data (Figure LABEL:VENDI_countries_diversity) and video data (Figure LABEL:VENDI_video_diversity), showcasing the applicability of the metric across different domains.
@@
We observe in Figure LABEL:VENDI_stylegan3_convergence that the convergence behavior is illustrated across various values of $\psi$. The results indicate that, for a fixed bandwidth $\sigma$, the truncated, Nyström, and FKEA variants of the VENDI score converge to the truncated VENDI statistic. As demonstrated in Figure LABEL:VENDI_ffhq_truncation of the main text, this truncated VENDI statistic effectively captures the diversity characteristics inherent in the underlying dataset.
We note that in presence of incremental changes to the diversity of the dataset, finite-dimensional kernels, such as cosine similarity kernel, remain relatively constant. This effect is illustrated in Figure LABEL:VENDI_styleganxl_convergence, where increase in truncation factor $\psi$ results in incremental change in diversity. This is one of the cases where infinite-dimensional kernel maps with a sensitivity (bandwidth) parameter $\sigma$ are useful in controlling how responsive the method should be to the change in diversity.
B.1 Approximation error of VENDI scores
In this section, we demonstrate the numerical stability of all VENDI score variants. Figure LABEL:error_bounds presents the standard deviation (expressed as a percentage) across varying sample sizes. The results indicate that, for each sample size $n$, the metric exhibits relatively low variance when computed across multiple sample sets drawn from the same underlying distribution, highlighting the robustness of the VENDI score computation. We present the results for ImageNet and FFHQ datasets.
B.2 Bandwidth $\sigma$ Selection

### v3

@@ A.5 Proof of Theorem 3
Proof of Part (a). As defined by Ospanov et al. [2024b], the FKEA method uses the eigenvalues of $t$ random Fourier frequencies $\omega_{1},\ldots,\omega_{t}$ where for each $\omega_{i}$ they consider two features $\cos(\omega_{i}^{\top}x)$ and $\sin(\omega_{i}^{\top}x)$. Following the definitions, it can be seen that $k(x,x^{\prime})=\mathbb{E}_{\omega\sim p_{\omega}}\bigl{[}\cos(\omega^{\top}(x-x^{\prime}))\bigr{]}$ which is approximated by FKEA as $\frac{1}{t}\sum_{i=1}^{t}\cos(\omega_{i}^{\top}(x-x^{\prime}))$. Therefore, if we define kernel matrix $K_{i}$ as the kernel matrix for $k_{i}(x,x^{\prime})=\cos(\omega_{i}^{\top}(x-x^{\prime}))$, then we will have
@@
On the other hand, we note that $\|\frac{1}{n}K\|_{\mathrm{HS}}\leq 1$ holds as the kernel function is normalized and hence $|k(x,x^{\prime})|\leq 1$. Since the Frobenius norm is the $\ell_{2}$-norm of the vectorized version of the matrix, we can apply Vector Bernstein inequality in Lemma 1 to show that for every $0\leq\epsilon\leq 2$:
@@
Then, we apply the Hoffman-Wielandt inequality to show that for the sorted eigenvalue vectors of $\frac{1}{n}K$ (denoted by $\widehat{\boldsymbol{\lambda}}_{n}$) and $\frac{1}{n}K^{\mathrm{FKEA}(t)}$ (denoted by ${\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}$) we will have $\|\widehat{\boldsymbol{\lambda}}_{n}-{\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}\|_{2}\leq\|\frac{1}{n}K^{\mathrm{FKEA}(t)}-\frac{1}{n}K\|_{\mathrm{HS}}$, which together with the previous inequality leads to
@@
Proof of Part (b). To show this theorem, we use Theorem 3 from [Xu et al., 2015], which shows that if the $r$th largest eigenvalue of the kernel matrix $\frac{1}{n}K$ satisfies $\lambda_{r}\leq\frac{\tau}{n}$, then given $t\geq Cr\log(n)$ ($C$ is a universal constant), the following spectral norm bound will hold with probability $1-\frac{2}{n^{3}}$:
@@ Appendix B Additional Numerical Results
In this section, we present supplementary results concerning the evaluation of diversity and the convergence behavior of different variants of the Vendi score. We extend the convergence experiments discussed in the main text to include the truncated StyleGAN3-t FFHQ dataset (Figure 12) and the StyleGAN-XL ImageNet dataset (Figure 13). Furthermore, we demonstrate that the truncated Vendi statistic effectively captures the diversity characteristics across various data modalities. Specifically, we conducted similar experiments as shown in Figures 7 and 6 on text data (Figure 9) and video data (Figure 11), showcasing the applicability of the metric across different domains.
Figure 9: Diversity evaluation of Vendi scores on synthetic text dataset about 100 countries generated by GPT-4 with varying number of countries. The setting is based on text-embedding-3-large embedding and bandwidth $\sigma=0.5$
@@
Figure 11: Diversity evaluation of Vendi scores on Kinetics400 dataset with varying number of classes. The setting is based on I3D embedding and bandwidth $\sigma=4.0$
We observe in Figure 12 that the convergence behavior is illustrated across various values of $\psi$. The results indicate that, for a fixed bandwidth $\sigma$, the truncated, Nyström, and FKEA variants of the Vendi score converge to the truncated Vendi statistic. As demonstrated in Figure 6 of the main text, this truncated Vendi statistic effectively captures the diversity characteristics inherent in the underlying dataset.
We note that in presence of incremental changes to the diversity of the dataset, finite-dimensional kernels, such as cosine similarity kernel, remain relatively constant. This effect is illustrated in Figure 13, where increase in truncation factor $\psi$ results in incremental change in diversity. This is one of the cases where infinite-dimensional kernel maps with a sensitivity (bandwidth) parameter $\sigma$ are useful in controlling how responsive the method should be to the change in diversity.
Figure 12: Statistical convergence of Vendi score for different sample sizes on StyleGAN3 generated FFHQ data at various truncation factors $\psi$: (Left plot) finite-dimension cosine similarity kernel (Right plot) infinite dimension Gaussian kernel with bandwidth $\sigma=35$. DinoV2 embedding (dimension 768) is used in computing the scores.
Figure 13: Statistical convergence of Vendi score for different sample sizes on StyleGAN-XL generated ImageNet data: (Left plot) finite-dimension cosine similarity kernel (Right plot) infinite dimension Gaussian kernel with bandwidth $\sigma=40$. DinoV2 embedding (dimension 768) is used in computing the scores.
Table 1: Statistical convergence of diversity scores for different sample size on DALL-E 3 generated MSCOCO data
$n$ VENDI-1.0 RKE Vendi-t FKEA-Vendi Nystrom-Vendi Recall Coverage
2000 239.91 13.47 239.91 228.69 239.91 0.76 0.86
4000 315.35 13.51 315.35 280.68 315.35 0.81 0.87
6000 357.15 13.56 346.27 310.9 345.49 0.83 0.91
8000 392.36 13.56 354.8 329.56 357.41 0.87 0.91
Table 2: Statistical convergence of diversity scores for different sample size on SDXL generated MSCOCO data
$n$ VENDI-1.0 RKE Vendi-t FKEA-Vendi Nystrom-Vendi Recall Coverage
2000 187.17 10.65 187.17 173.06 187.18 0.78 0.85
4000 236.49 10.7 236.49 222.78 236.08 0.82 0.87
6000 264.82 10.7 258.21 236.37 257.34 0.86 0.87
8000 289.08 10.71 265.84 251.59 266.23 0.86 0.86
10000 304.44 10.72 267.39 256.24 268.34 0.86 0.87
Table 3: Compilation time (in seconds) of different Vendi scores with increasing sample size
Metric samples $n$
10000 20000 30000 40000 50000 60000 70000
Vendi 97s 631s 1868s - - - -
FKEA-Vendi 19s 36s 53s 71s 88s 105s 124s
Nystrom-Vendi 31s 44s 78s 91s 112s 136s 164s
B.1 Bandwidth $\sigma$ Selection
@@
Appendix C Selection of Embedding space
To show that proposed truncated Vendi score remains feasible under arbitrary embedding selection, we conducted experiments from Figures 6 and 7. Figures 14, 15, 16 and 17 extend the results to CLIP Radford et al. [2021] and SWaV Caron et al. [2020] embeddings. These experiments demonstrate that FKEA, Nyström and $t$-truncated Vendi correlate with increasing diversity of the evaluated dataset. We emphasize that proposed statistic remains feasible under arbitrary embedding space that is capable of mapping image samples into a latent space.
Figure 14: Diversity evaluation of Vendi scores on ImageNet dataset with varying number of classes based on CLIP embedding and bandwidth $\sigma=5.0$
Figure 15: Diversity evaluation of Vendi scores on truncated StyleGAN3 generated FFHQ with varying truncation coefficient $\psi$ based on CLIP embedding and bandwidth $\sigma=5.0$
Figure 16: Diversity evaluation of Vendi scores on ImageNet dataset with varying number of classes based on SWaV embedding and bandwidth $\sigma=1.0$
Figure 17: Diversity evaluation of Vendi scores on truncated StyleGAN3 generated FFHQ with varying truncation coefficient $\psi$ based on SwAV embedding and bandwidth $\sigma=1.0$